OpenStudy (anonymous):

the sum of the measures of two angles is 180. One angle measures 3a and the other angle measures 2a-25 Find a. Then find the measure of each angle.

OpenStudy (anonymous):

3a + (2a - 25) = 180 ...........5a -25 = 180 .................5a = 180+25 .................5a = 205 ...................a = 41 angle with measure 3a: 3a = 3(41) = 123 angle with measure 2a - 25: 2a - 25 = 2(41) - 25 = 82 - 25 = 57